jdog1882:

Mike is working on solving the exponential equation 37x = 12; however, he is not quite sure where to start. Using complete sentences, describe to Mike how to solve this equation. Hint: Use the change of base formula: log base b of y equals log y over log b.

ZombieLover:

What would you choose?

shoooootooootodoroki:

Mike should multiply 37 and x and it should equal 12

Mercury:

This is an exponential equation, which means it's most likely 37^x = 12 rather than 37x = 12 in this case, you use logarithms to change the base \[\log_{b}y=\frac{ \log(y) }{ \log(b) }\] like the problem states. you can take \[\log_{37}\]on both sides. from there, you can isolate x on the left side by cancelling out log_37)(37^x) and be left with log_37 (12) on the right side. use change of base to simplify the right side.